OpenStudy (anonymous):

p(t)=under root t-4/3t-21 find the domain

OpenStudy (anonymous):

the first one is the correct one

OpenStudy (anonymous):

he unde root is just on the top

OpenStudy (anonymous):

the under root is just on the top

OpenStudy (phi):

\[p(t)= \frac{ \sqrt{t-4} }{ 3(t-7) }\]

OpenStudy (anonymous):

yes

OpenStudy (phi):

First, let's do the no divide by zero. What bad t do we exclude?

OpenStudy (anonymous):

the one on the top

OpenStudy (anonymous):

?

OpenStudy (phi):

You have a fraction. If the bottom is zero, that is bad. what t must we not allow? (otherwise we will divide by 0)

OpenStudy (anonymous):

so on the top t>4

OpenStudy (phi):

so on the top t>4 ALMOST. t >= 4 RIGHT?

OpenStudy (phi):

sqrt(0) is 0 that is ok (it is defined)

OpenStudy (phi):

now find the bad t that makes the bottom zero.

OpenStudy (anonymous):

yes right and what abut 3 and 7?

OpenStudy (anonymous):

that is 7 and 3

OpenStudy (anonymous):

bad one which makes t 0

OpenStudy (anonymous):

or just -7

OpenStudy (phi):

to find the bad t: solve 3(t-7)=0

OpenStudy (phi):

divide by 3 add 7

OpenStudy (anonymous):

3 divided by 7?

OpenStudy (phi):

3(t-7)=0 divide both sides by 3 (t-7)= 0/3 or t-7=0 can you finish?

OpenStudy (anonymous):

t=-7

OpenStudy (anonymous):

I mean t=7

OpenStudy (phi):

I guess not!

OpenStudy (phi):

t=7, much better. so you have all real numbers t where t>=4 and t≠7 as the domain.

OpenStudy (anonymous):

why t is not equal to 7?

OpenStudy (anonymous):

why not t>7 .

OpenStudy (phi):

The t=7 makes the bottom 0. NO DIVIDE BY ZERO.

OpenStudy (phi):

For the top to be ok t must be 4 or bigger. But when you get to 7, you have to toss it, because the bottom becomes 0 when t=7.

OpenStudy (anonymous):

ok

OpenStudy (phi):

is that a ok (I don't get it) or an ok (I understand) ?

OpenStudy (anonymous):

hm well I still did not get how t=7 is not equal to 7:(

OpenStudy (anonymous):

I did got the concept of how to find the domain. How to find the good and bad x .

OpenStudy (anonymous):

I also learnt tht we have to simplfy and then see if x> any number.

OpenStudy (phi):

The domain is all the "good t's" when you found t=7 makes the denominator 0, you found a "bad t" so you take the good t's from the top, and then say, by the way, don't include this bad t. You say t ≠7 (read this as "t is not allowed to be 7")

OpenStudy (anonymous):

ohhhh . I was looking it from another angle. I was thinkinh of the 4 on the top..

OpenStudy (anonymous):

t=7/0 not ok

OpenStudy (anonymous):

right?

OpenStudy (phi):

Let's simplify this: what is the domain for f(x)= 1/(x-7) ? It is all real numbers , except 7 you say x is all real numbers, x≠7 For this problem, it is almost the same, except x is all real numbers ≥ 4, except for 7 x is all real numbers ≥4, x≠7

OpenStudy (anonymous):

where the 4 came?

OpenStudy (phi):

you have \[ \sqrt{t-4} \] in the numerator. You want the stuff inside the root to be 0 or positive. (no negatives) so you want t-4≥0 or t ≥4

OpenStudy (anonymous):

ok yes this part is under stood.

OpenStudy (anonymous):

I am really worried how will I pass this class?

OpenStudy (phi):

You seem to be learning it. I'll be back tomorrow about the same time. Meanwhile others will help you.

OpenStudy (anonymous):

thanx alot. I wil take a break too.